# Fundamental computational limits of weak learnability in high-dimensional multi-index models

### v2

@@ Appendix A Optimality of AMP
Note that the state evolution Equation 7 is closely related to the $\sup\inf$ problem in Equation 29. Indeed, remarking that the update function $F$ in Equation 7 is precisely the gradient of the entropy $H_{Y}$ in Equation 8, one can show that state evolution is equivalent to gradient descent in the objective defined by Equation 29 [Aubin et al., 2019]. This non-trivial fact implies that whenever Equation 29 has a single minima, AMP optimally estimates the posterior marginals.
Finally, note that by construction $\hat{{{\bm{{W}}}}}_{\rm bo}$ is the optimal estimator of ${{\bm{{W}}}}^{\star}$ given the data $(X,y)$ (in the MMSE sense). Therefore, the rank of $M^{\star}$ defines the dimension of the statistically optimal subspace reconstruction at sample complexity $\alpha\coloneqq\nicefrac{{n}}{{d}}$. The fact that AMP follows state evolution is proven for such problems in [Javanmard and Montanari, 2013, Gerbelot and Berthier, 2023].
Appendix B Proofs of the main results
B.1 Linear Approximation
@@
$\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\,\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})^{\top}\right]\,,$ (31)
@@
B.2 Proof of Lemma 4.1
@@
Thus, $\mathcal{F}(M)$ is a cone-preserving linear map. The generalised Perron-Frobenius theorem for cone-preserving maps [Krein and Rutman, 1948] then implies that the operator $\mathcal{F}({{\bm{M}}})$ admits a unique eigenvector $M^{*}\in\mathcal{S}^{+}_{p}$ corresponding to the largest eigenvalue $\lambda_{\mathcal{F}}$ such that for any $M\in\mathcal{S}^{+}_{p}\perp M^{*}$:
$\mathcal{F}({{\bm{M}}})<\lambda_{\mathcal{F}}\norm{{{\bm{M}}}}_{F}.$ (34)
Furthermore, all other eigenvalues are strictly smaller than $\lambda_{\mathcal{F}}$. Subsequently, Lemma B.1 implies that $F({{\bm{M}}})$ is stable at ${{\bm{M}}}=0$ if and only if $\alpha\leq\frac{1}{\lambda_{\mathcal{F}}}=\alpha_{c}$.
B.3 Proof of Theorem 4.2
Lemma 2.1 allows us to map the behavior of the variable ${{\bm{M}}}^{t}$ in state-evolution to high-probability statements for the limiting overlaps Suppose that $A\in\mathcal{S}^{+}_{p}$ is a full rank matrix as in Theorem 4.2. Applying Lemmas B.1 and 3.2, we obtain that for ${{\bm{M}}}^{0}=\epsilon A$:
$\norm{{{\bm{M}}}^{t+1}}_{F}\leq\alpha\lambda_{\mathcal{F}}\norm{{{\bm{M}}}^{t}}_{F}+\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F},$ (35)
for some constant $C$. Now, suppose that $\alpha<\frac{1}{\lambda_{F}}$ and $\epsilon<\frac{1}{C}(1-\alpha\lambda_{\mathcal{F}}-\kappa)$, for some $0<\kappa<1-\alpha\lambda_{\mathcal{F}}$ then inductively, we obtain:
$\norm{{{\bm{M}}}^{t+1}}_{F}<(1-\kappa)\norm{{{\bm{M}}}^{t}}_{F},$ (36)
for all $t>0$. Implying that $\norm{{{\bm{M}}}^{t}}_{F}\rightarrow 0$.
Now, suppose that $\alpha>\alpha_{c}$ or equivalently $\alpha\lambda_{F}>1$. Since $A$ is full-rank admits a decomposition:
$\operatorname{tr}(A,{{\bm{M}}}^{\star})=\theta$ (37)
for some $\theta>0$. From Lemma 4.1 and B.1, we obtain:
$\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})\geq\alpha\lambda_{\mathcal{F}}\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (38)
For $\norm{{{\bm{M}}}^{t}}\leq\frac{\theta}{C}(\alpha\lambda_{\mathcal{F}}-1-\kappa)$ for some $\kappa>0$, we inductively obtain:
@@
implying that $\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})$ grows as $\omega(e^{\kappa t})$ for sufficiently small $\epsilon$.
@@
$\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v=\mathbb{E}_{y}\left[(\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p})\mathbf{v})^{2}\right]>0,$ (40)
while for any $A\in\mathcal{S}^{+}_{p}$:
$\mathbf{v}^{\top}\mathcal{F}(A)v\geq 0,$ (41)
since $\mathcal{F}(A)\in\mathcal{S}^{+}_{p}$. Define:
$\nu_{\mathcal{F}}=\inf_{\mathbf{v}\in\mathbb{E},\norm{v}=1}\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})v.$ (42)
Since $\mathbf{v}$ lies in a compact set, $\nu_{\mathcal{F}}>0$. Let $A^{\star}$ denote the projection of $A$ along $E^{\star}$ and consider its eigendecomposition $A^{\star}$:
$A^{\star}=\sum_{i=1}^{p}\lambda_{i}\mathbf{v}_{i}\mathbf{v}^{\top}_{i}$ (43)
We
$\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t+1})\mathbf{v}_{i}\geq\alpha\nu_{\mathcal{F}}\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t})\mathbf{v}_{i}-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (44)
@@
B.5 Proof of Theorem 5.1
By assumption, $M_{U}$ is a fixed point of $F(M)$. Therefore, Equation 7 implies that ${\bm{g}}_{\rm out}(y,{\bm{\omega}},{\bm{I}}-\sqrt{M}_{U})$ almost surely lies in $U$ and thus $T^{\star}_{U}$ is empty. The proof of Theorem 5.1 then follows that of 4.2 by considering the following linearized operator along $U^{\perp}$

### v3

@@ Appendix A Bayes optimality of AMP
Note that the state evolution Equation 9 is closely related to the $\sup\inf$ problem in Equation 29. Indeed, remarking that the update function $F$ in Equation 9 is precisely the gradient of the entropy $H_{Y}$ in Equation 10, one can show that state evolution is equivalent to gradient descent in the objective defined by Equation 29 [Aubin et al., 2019]. This non-trivial fact implies that whenever Equation 29 has a single minima, AMP optimally estimates the posterior marginals, at least if its state evolution converges to this fixed point.
Finally, note that by construction $\hat{{{\bm{{W}}}}}_{\rm bo}$ is the optimal estimator of ${{\bm{{W}}}}^{\star}$ given the data $(X,y)$ (in the MMSE sense). Therefore, the rank of $M^{\star}$ defines the dimension of the statistically optimal subspace reconstruction at sample complexity $\alpha\coloneqq\nicefrac{{n}}{{d}}$. The fact that AMP follows state evolution is proven for such problems in [Javanmard and Montanari, 2013, Gerbelot and Berthier, 2023]. Specifically, we refer to the example 4.4 in Gerbelot and Berthier [2023] on Matrix-valued random variables.
Appendix B Bayes-AMP in the presence of side-information
The algorithm is easily generalized to the case with side information of the type eq.(18):
${\bf S}=\sqrt{\lambda}W^{\star}+\sqrt{1-\lambda}Z,$ (30)
The effect of the additional side-information is equivalent to a modification of the prior on $W$ to $P^{S}_{W}(W)$, defined as:
$P^{S}_{W}({{\bm{{W}}}})$ $\propto P({\bm{S}}|{{\bm{{W}}}})P_{W}({{\bm{{W}}}})={\cal N}(\sqrt{\lambda}{\bf S},\sqrt{1-\lambda}\mathbf{I})$
This is a simple consequence of Bayes theorem:
$P({{\bm{{W}}}}|{\bm{y}},{\bm{X}},{\bm{S}})\propto P({\bm{y}},{\bm{X}},{\bm{S}}|{{\bm{{W}}}})P_{W}({{\bm{{W}}}})=P({\bm{y}},{\bm{X}}|{{\bm{{W}}}})P({\bm{S}}|{{\bm{{W}}}})P_{W}({{\bm{{W}}}})$ (31)
In turns, this change updates of $\hat{{{\bm{{W}}}}}$ and $\hat{{\bm{C}}}$ in Algorithm 1 as described in Aubin et al. [2019]:
$\hat{{{\bm{{W}}}}}_{k}^{t+1}=({\bm{I}}_{p}(1-\lambda)+{\bm{A}}_{k}^{t})^{-1}(\bm{b}_{k}^{t}(1-\lambda)+\sqrt{\lambda}{\bm{A}}_{k}^{t}{\bm{S}})\,,\qquad\hat{{\bm{C}}}_{k}^{t+1}=({\bm{I}}_{p}(1-\lambda)+{\bm{A}}_{k}^{t})^{-1}\,.$ (32)
It is easily checked that the limit $\lambda\to 0$ gives back the equations in Algorithm 1. Lemma 2.1 is then modified with $G(X)$ in equation (9) being replaced by
$G(X)=(X(1-\lambda)+\lambda)(1+X(1-\lambda))^{-1}\,.$ (33)
Appendix C Trivially-coupled Subspaces
In this section, we discuss and illustrate examples of $g$ where $T^{\star}_{U}\neq\emptyset$ for a learned subspace $U$. Our prototypical example throughout the discussion will be:
$g(z_{1},z_{2},z_{3})=z^{2}_{1}+z^{2}_{2}+\text{sign}(z_{1}z_{2}z_{3}).$ (34)
Let $e_{1},e_{2},e_{3}$ denote the directions corresponding to components $z_{1},z_{2},z_{3}$ respectively. It is easy to check that for the above example, $E^{\star}=\{e_{1},e_{2}\}$ while $H^{\star}_{E}=\{e_{3}\}$. Therefore, by Theorem 4.2, for large enough $\alpha$ the ${{\bm{M}}}^{t}\geq\delta M_{E}^{\star}$ where $M_{E}^{\star}$ spans $z_{1},z_{2}$ after $t=\mathcal{O}(\log\frac{1}{\lambda})$ iterations. Subsequently, since $T^{\star}_{E^{\star}}=\{e_{3}\}$, it gains an overlap along $e_{3}$ (independent of $\lambda$). Unlike sequential learning of directions in $E^{\star}_{U}$, however, the time-steps for recovery of $\{e_{1},e_{2}\}$ and $\{e_{3}\}$ (upto a threshold independent of $\lambda$) different only by a constant. The recovery of $e_{3}$ can therefore be interpreted as instantaneous w.r.t $\{e_{1},e_{2}\}$.
Appendix D Proofs of the main results
In this section, we describe the proofs of our main results, starting with a Trivial proof of the learning of Trivial subspaces (Theorem 3.2) and the form of linearization that will be utilized throughout the analysis.
D.1 Theorem 3.2 and Linearization
Lemma 3.1 and Theorem 3.2 are direct consequences of the following observation:
Proposition D.1.
Let $\mathbf{v}\in\mathbb{R}^{p}$ be arbitrary. Then, starting from ${{\bm{M}}}^{0}=0$, $\mathbf{v}^{\top}{{\bm{M}}}^{1}\mathbf{v}>0$ if and only if:
$\alpha\mathbb{E}\left[\mathbf{v}^{\top}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right){\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right)^{\top}\mathbf{v}\right]>0$ (35)
Proof.
The above proposition follows directly from the form of $F$ in Lemma 2.1 and the observation that the mapping $G(X)$ preserves the span of $X$. ∎
The expression in Equation 35 can be rewritten as:
$\mathbb{E}\left[(\mathbf{v}^{\top}{\bm{g}}_{\rm out}\!\left(Y^{t},0,{\bm{I}}_{p}\right))^{2}\right]$ (36)
Since the operand inside the expectation is non-negative, the expectation is non-zero if and only if the operand vanishes identically. We therefore obtain Lemma 3.1 and Theorem 3.2.
D.1.1 Linearization without side-information
@@
D.1.2 Linearization in the presence of side-information
Analogous to Lemma D.2, we can linearize state-evolution in the presence of side-information in the limit of small $M,\lambda$.
Lemma D.3.
Let $\mathcal{F}(\delta{{\bm{M}}})$ be a linear operator on the cone $\mathcal{S}^{+}_{p}$ of PSD matrices of dimension $p$, defined by:
$\mathcal{F}({{\bm{M}}})\mathrel{\mathop{:}}=\,\mathbb{E}_{y}\left[\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})\,{{\bm{M}}}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p}-{{\bm{M}}})^{\top}\right]$ (39)
Then the following approximation holds:
$F({{\bm{M}}})\approx\alpha\mathcal{F}(\delta{{\bm{M}}})+\sqrt{\lambda}\mathbf{I}_{d}+\mathcal{O}(\alpha\norm{\delta{{\bm{M}}}}_{F}^{2})+\mathcal{O}(\lambda),$ (40)
Proof.
The above linearization is a direct consequence of Lemma D.2 along with the updated form of $G(X)$ in Equation 33. ∎
D.2 Proof of Lemma 3.3
Proof.
This is a direct consequence of the Tower law of expectation: $\mathbb{E}\left[\mathcal{T}(y)\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle\right]=\mathbb{E}_{y}\left[\mathcal{T}(y)\mathbb{E}_{\mathbf{x}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]\right]$ The statement then follows by noting that ${\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})^{\top}\mathbf{v}=\lim_{d\rightarrow\infty}\mathbb{E}_{{\bm{x}}}\left[\langle\mathbf{v}^{\top}W^{\star},\mathbf{x}\rangle|y\right]$. The converse for any $v\in T^{\star}$ then follows by setting $\mathcal{T}(y)={\bm{g}}_{\rm out}(y,\mathbf{0},{\bm{I}}_{p})$.∎
D.3 Proof of Lemma 4.1
@@ D.1.2 Linearization in the presence of side-information
Thus, $\mathcal{F}(M)$ is a cone-preserving linear map. The generalised Perron-Frobenius theorem for cone-preserving maps [Krein and Rutman, 1948] then implies that the operator $\mathcal{F}({{\bm{M}}})$ admits a unique eigenvector $M^{*}\in\mathcal{S}^{+}_{p}$ corresponding to the largest eigenvalue $\nu_{\mathcal{F}}$ such that for any $M\in\mathcal{S}^{+}_{p}\perp M^{*}$:
$\mathcal{F}({{\bm{M}}})<\nu_{\mathcal{F}}\norm{{{\bm{M}}}}_{F}.$ (43)
Furthermore, all other eigenvalues are strictly smaller than $\lambda_{\mathcal{F}}$. Subsequently, Lemma D.2 implies that $F({{\bm{M}}})$ is stable at ${{\bm{M}}}=0$ if and only if $\alpha\leq\frac{1}{\nu_{\mathcal{F}}}=\alpha_{c}$.
D.4 Proof of Theorem 4.2
Lemma 2.1 allows us to map the behavior of the variable ${{\bm{M}}}^{t}$ in state-evolution to high-probability statements for the limiting overlaps.
Applying Lemmas D.3 and 3.3, we obtain that for small enough ${{\bm{M}}}^{t+1},\lambda$:
$\norm{{{\bm{M}}}^{t+1}}_{F}\leq\alpha\nu_{\mathcal{F}}\norm{{{\bm{M}}}^{t}}_{F}+\lambda+C_{1}\norm{{{\bm{M}}}^{t}}^{2}_{F}+C_{2}\norm{{{\bm{M}}}^{t}}^{2}_{F}$ (44)
for some constants $C_{1},C_{2}$. Now, suppose that $\alpha<\frac{1}{\nu_{F}}$. Then for $\norm{{{\bm{M}}}^{t}}_{F}\geq\frac{1}{(1-\alpha\nu_{\mathcal{F}})}\lambda$, we have:
$\alpha\nu_{\mathcal{F}}\norm{{{\bm{M}}}^{t}}_{F}+\lambda\leq\norm{{{\bm{M}}}^{t}}_{F}.$ (45)
For small enough $\lambda$, we have that:
${{\bm{M}}}^{0}\approx\lambda<\frac{1}{(1-\alpha\nu_{\mathcal{F}})}\lambda.$ (46)
Subsequently, by induction, we obtain that:
$\sup_{t\geq p}\norm{{{\bm{M}}}^{t+1}}_{F}\leq\frac{1}{(1-\alpha\nu_{\mathcal{F}})}\lambda$ (47)
for all $t>0$. This proves the first part of the Theorem.
Now, suppose that $\alpha>\alpha_{c}$ or equivalently $\alpha\nu_{F}>1$.
From Lemma 4.1 and D.3, we obtain:
$\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})\geq\alpha\nu_{\mathcal{F}}\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})+\lambda-C_{1}\norm{{{\bm{M}}}^{t}}^{2}_{F}-C_{2}\lambda^{2},$ (48)
for some constants $C_{1},C_{2}>0$.
Since $\alpha\nu_{F}>1$, we obtain that $\operatorname{tr}({{\bm{M}}}^{1},{{\bm{M}}}^{\star})>\lambda>0$ for small enough $\lambda$.
Let $10<\kappa<\nu_{F}-$ 1 be fixed. Equation 48 implies that for $\frac{\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})}{\norm{{{\bm{M}}}^{t}}^{2}_{F}}<\frac{\alpha\nu_{F}-1-\kappa}{C_{1}}$, and small enough $\lambda$:
@@
implying that $\operatorname{tr}({{\bm{M}}}^{t+1},{{\bm{M}}}^{\star})$ grows as $\omega(e^{\kappa t})$. Since at $t=1$, $\frac{\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})}{\norm{{{\bm{M}}}^{t}}^{2}_{F}}=\mathcal{O}(\frac{1}{\lambda})$, Equation 44 implies that the condition $\frac{\operatorname{tr}({{\bm{M}}}^{t},{{\bm{M}}}^{\star})}{\norm{{{\bm{M}}}^{t}}^{2}_{F}}<\frac{\alpha\nu_{F}-1-\kappa}{C_{1}}$ holds for time $\mathcal{O}(\log\frac{1}{\lambda})$, which along with Equation 49 allows recovery upto an overlap $\delta$ independent of $\lambda$.
@@
$\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})\mathbf{v}=\mathbb{E}_{y}\left[(\mathbf{v}^{\top}\partial_{{\bm{\omega}}}{\bm{g}}_{\rm out}(y,0,{\bm{I}}_{p})\mathbf{v})^{2}\right]>0,$ (50)
Define the smallest eigenvalue of $\mathcal{F}$:
$\nu^{\star}_{\mathcal{F}}=\inf_{\mathbf{v}\in E^{\star},\norm{v}=1}\mathbf{v}^{\top}\mathcal{F}(\mathbf{v}\mathbf{v}^{\top})\mathbf{v}.$ (51)
Since $\mathbf{v}$ lies in a compact set, $\nu^{\star}_{\mathcal{F}}>0$.
Let $\mathbf{v}_{1},\cdots,\mathbf{v}_{k}$ denote an orthonormal basis for $E^{\star}$
$\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t+1})\mathbf{v}_{i}\geq\alpha\nu_{\mathcal{F}}\mathbf{v}^{\top}_{i}({{\bm{M}}}^{t})\mathbf{v}_{i}+\lambda-\alpha C\norm{{{\bm{M}}}^{t}}^{2}_{F}.$ (52)
@@
D.6 Staircase Dynamics and the proof of Proposition 5.1
Before proving Proposition 5.1, we start by proving a monotonicity property of the AMP iterates, which in-turn implies that the AMP iterates gaining overlap along a certain subspace must eventually reach an approximate fixed point.
D.6.1 Monotonicity of the AMP Iterates
We begin by showing that the overlaps of the Bayes-AMP iterates is non-decreasing in the following sense:
Lemma D.4.
Consider the AMP algorithm with the Bayes-optimal choice of $f_{t},g_{t}$. Then for any $t>0$:
${{\bm{M}}}^{(t+1)}\succeq{{\bm{M}}}^{(t)}$ (56)
Proof.
Our proof relies on a characterization of the state-evolution for the Bayes-AMP iterates as the solution to certain low-dimensional inference problems. See for instance Barbier et al. [2019], Aubin et al. [2019]. Concretely, consider the following sequence of denoising problems over variables $Z^{\star}\in\mathbb{R}^{k},W^{\star}\in\mathbb{R}^{k}$:
$Y_{t}=g(\sqrt{{{\bm{M}}}^{(t)}}\xi+\sqrt{\mathbb{I}_{k}-{{\bm{M}}}^{(t)}}Z^{\star})$
$\tilde{Y}_{t}=\sqrt{\tilde{M}}_{t}W^{\star}+Z$
where $Z,\xi\sim\mathcal{N}(0,\mathbf{I}_{k})$ and $\tilde{M}_{t}$ is obtained as: